g#/assignment/6338912 VitemView?offset- next&assignmentProblemID-100453845 80 g of Na2SO4 in 270 mL, of Na2SO4 solution Express your answer using two significant figures. Submit Request Answer ? Part B 44 g of sucrose in 300 mL of a carbonated drink Express your answer using two significant figures. %m/v
Answer #1

Part a

(m/V) = mass of solute / volume of solution

80 g of Na2SO4 dissolving into 270 mL of Na2SO4

Mass/volume % = 80*100/270

= 29.63%

Part b

% m/V = 44*100/300

= 14.66%